OpenStudy (cggurumanjunath):

FIND VELOCITY,SPEED AND ACCELERATION AND DESCRIBE MOTION OF PARTICLE WHOSE POSITION AT TIME t is given by the equation below.

OpenStudy (cggurumanjunath):

\[r(t)=3\cos (wt)i+4\cos (wt)j+5\sin(wt)k\]